OpenStudy (anonymous):

Solve the system 2x – 3y = 7 and 8x – 12y = 5 by using graph paper or graphing technology. What is the solution to the system?

OpenStudy (p0sitr0n):

isolate a variable and subsitute its equation in the other equation

OpenStudy (p0sitr0n):

2x -3y =7 2x = 3y +7 x =(3y+7)/2 \[8((3y+7)/2) - 12y =5\]